Which company has the greater return on equity based on the following information: Company 1 2...

  1. Which company has the greater return on equity based on the following information:

Company

1

2 – Capital

Revenue

          35,000

       39,000

Other expenses

          12,000

       12,500

Interest

922

838

Lease categorization

Operating

>75 of useful life

Lease payment

            5,750

5,223

Life of lease

                   4

                5

Current year of lease

                   2

                3

Total equity

            500,000

         620,000

Incremental borrowing rate

6%

6%
Answer #1

Company B is profitable in the current year :

Particulars Company 1 Company 2
Revenue           35,000           39,000
Other Expenses         (12,000)         (12,500)
Interest           (1,195)           (1,086)
ROU Depreciation           (4,981)           (4,400)
          16,823           21,014
Capital        500,000        620,000
Rate of Return 3.36% 3.39%

Working Notes :

For company B it is mentioned that lease categorization is <75% of useful life. As per IFRS and US GAAP where lease asset is having lease period less than 75% of useful life is not a finance lease so it will also be categorized as operating lease.

Under operating lease lessee accounts for :

Deprecation on Right to Use(ROU) Asset and also accounts for interest expense.

The ROU = Net present value of lease payment

Company A = NPV(6%,5750,5750, 5750,5750,) = 19,924, ROU Depreciation 4981 = 19,924/4

Company B = NPV(6%,5223,5223,5223,5223,5223) = 22,001, ROU Depreciation 4,400 = 22,001/5

Company A calculation of Implied Interest :

Opening Balance Interest Lease Payment Principal repayment Closing balance of principal
$19,924 $1,195 $5,750 $4,555 $15,370
$15,370 $922 $5,750 $4,828 $10,542
$10,542 $633 $5,750 $5,117 $5,425
$5,425 $325 $5,750 $5,425 $0

Company B calculation of Implied Interest :

Opening Balance Interest Lease Payment Principal repayment Closing balance of principal
$22,001 $1,320 $5,223 $3,903 $18,098
$18,098 $1,086 $5,223 $4,137 $13,961
$13,961 $838 $5,223 $4,385 $9,576
$9,576 $575 $5,223 $4,648 $4,927
$4,927 $296 $5,223 $4,927 ($0)